Find the pH of a 0.130 M solution of a weak monoprotic acid having Ka= 1.1×10−3.

Express your answer to two decimal places.
Answer #1

Dissociation equilibrium of monoprotic acid is

HA(aq) + H2O(l ) <------> H3O+(aq) + A-(aq)

Ka = [H3O+][A-]/[HA] = 1.1 ×10-3

at equilibrium

[HA] = 0.130 - x

[A-] = x

[H3O+] = x

so,

x2/( 0.130 - x) = 1.1 ×10-3

solving for x

x = 0.01142

[H3O+] = 0.01142M

pH = -log[H3O+]

pH = -log(0.01142M)

pH = 1.94